OpenStudy (anonymous):

De Moivres theorem (5-4i)^10 no need to solve this the question is if i'm expressing this as r^10(cos(10a)+isin(10a)) how do i find a? is it just tan-1(-4/5)? or do i have to think about the quadrants?

OpenStudy (anonymous):

you have to think about quadrants

OpenStudy (anonymous):

range of arctan is from -pi/2 to pi/2 so arctan will not give the right argument if you are in quadrant 2 or 3

OpenStudy (anonymous):

actually in this case you are ok because you are in quadrant 4 so you can just use arctan(-4/5)

OpenStudy (anonymous):

but 4th quadrant arguments ar expressed as - tan(|y/x|)^-1

OpenStudy (anonymous):

what if it was (-5-4i) and i was in quadrant 3?

OpenStudy (anonymous):

not sure what you mean. arctan(-4/5) is the number (angle) between -pi/2 and p/2 whose tangent if -4/5. this

OpenStudy (anonymous):

you are in quadrant three right? so if you take arctan(b/a) you get something in quadrant 1, directly across the circle. add pi to your answer in that case

OpenStudy (anonymous):

in third quadrant general argument representation is tan(|y/x)^-1 -pi

OpenStudy (anonymous):

i am user vicky007 is correct as well. there is no unique answer.

OpenStudy (anonymous):

@vicky your argument will be negative right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if it give the same result, should i just do tan-1(y/x) in each case and not bother with quadrants?

OpenStudy (anonymous):

@tsolrm the long and the short of it is since the range of arctan is restricted to -pi/2 to pi/2 you do have to check the quadrant.

OpenStudy (anonymous):

oh no not the same result!

OpenStudy (anonymous):

easy example take \[-1+i\]

OpenStudy (anonymous):

you are in quadrant 2. but if just use \[\tan^{-1}(\frac{b}{a})\] you get \[tan^{-1}(-1)=-\frac{\pi}{4}\]

OpenStudy (anonymous):

and that is definitely wrong.

OpenStudy (anonymous):

especially since \[\sqrt{2}(\cos(-\frac{\pi}{4})+i\sin(-\frac{\pi}{4}))=1-i\] not \[-1+i\]

OpenStudy (anonymous):

i've looked at past papers and they've never sticked a minus sign in the demoivres question it's always positive and therefore it's in quadrant 1. so i got used to doing tan-1(y/x) and that's it. so now i'm terrified that there will be a minus sign tomorrow

OpenStudy (anonymous):

but what you've said really cleared this up. thank you so much

OpenStudy (anonymous):

i am not sure what you mean by "it is alway positive" , but good luck, and yes make sure to know what quadrant you are in!

OpenStudy (anonymous):

i mean that they've always presented it as (something + i(something)) ^ something so the x and y have always been positive :)

OpenStudy (anonymous):

ooh well hope they do tomorrow too, but it is not such a big deal so long as you remember to check

OpenStudy (anonymous):

i'll need a miracle to pass this exam haha thanks again!